# Polynomial Expectation Property for Max-Polymatrix Games

### v1

@@ Abstract
We address an open question proposed by Papadimitriou and Roughgarden [1], which addresses the computability of correlated equilibria in a variant of polymatrix where each player’s utility is the maximum of their edge payoffs. We demonstrate that this max-variant game has the polynomial expectation property, and the results of [1] can thus be applied. We propose ideas for extending these findings to other variants of polymatrix games, as well as briefly address the broader question of necessity for the polynomial expectation property when computing correlated equilibria.
@@ 1.4 Polymatrix Games
In polymatrix games, the utility function $U$ can be reduced to a series of per-game utility functions $U_{pq}$, such that $U_{pq}(z,p,(s_{p},s_{q}))$ denotes player $p$’s utility in a two-player game against player $q$. In traditional polymatrix games, a player $p$’s utility is the total over all of their individual two-player games:
@@ 2 Polynomial Expectation Property for Max-Variant Polymatrix Games
By construction, $M$ is nonincreasing during the loop, so it suffices to subtract $x_{q}(k)$ from $c_{q}$ once $u_{q}(k)=M$ (i.e. $u_{q}(k)$ becomes the “current" utility). Note that, in the case of ties, we can assign the same (arbitrary) tiebreaking rule to the algorithm’s “argmax" and to the “$>$" operator, and this invariant will remain true.
@@
     $E(i)\leftarrow 0$triangleright exp. utility given fixed action $i$
     $c_{q}\leftarrow 1,\,\,\forall q\neq p$triangleright “CDF" variable for each player
@@
         $P\leftarrow x_{q^{*}}(0)\cdot\prod_{q\neq q^{*}}c_{q}$
         $E(i)\leftarrow E(i)+u_{q^{*}}(0)P$
@@
         Remove action 00 from $u_{q^{*}}$, shift other actions up
@@ 3.2 Necessity of Polynomial Expectation
The spirit of the original question proposed by Papadimitriou and Roughgarden was to address cases of games without polynomial expectation property, and argue whether they still have polynomial correlated equilibrium schemes. As it turns out, the example they provided, Max-Polymatrix, does indeed have the polynomial expectation property. This inspires the following general question: Papadimitriou and Roughgarden demonstrate that the polynomial expectation property is a sufficient condition for a polynomial correlated equilibrium scheme. To what extent is it necessary?
A small idea in this direction relates to the necessity of the computation of $UX^{\top}$ in the original algorithm in [1]. In particular, to solve the dual program $[UX^{\top}]\alpha\geq 0$, it is not entirely necessary to compute $UX^{\top}$ itself; it suffices to have a separation oracle which, given some input $\alpha$, returns either that it is feasible or a halfspace it violates (via the ellipsoid algorithm). This is equivalent to having an oracle which, given a distribution $X^{\top}\alpha$ [^1] Note that the entire vector $X^{\top}\alpha$ might be hard to compute, as we are linearly combining very large vectors. However, we can keep the representation as a linear combination of product distributions and treat “accessing” any element in this distribution as taking polynomial-time instead of constant., returns either that it is a correlated equilibrium or demonstrates a failed constraint. So, we have:
[…]

### v3

@@ Abstract
We address an open question proposed by Papadimitriou and Roughgarden [1] on the computability of correlated equilibria in a variant of polymatrix where each player’s utility is the maximum of their edge payoffs. We demonstrate that this max-variant game has the polynomial expectation property, and the results of [1] can thus be applied. We propose ideas for extending these findings to other variants of polymatrix games, as well as briefly address the broader question of necessity for the polynomial expectation property when computing correlated equilibria.
@@ 1.4 Polymatrix Games
In polymatrix games, the utility function $U$ can be reduced to a series of per-game utility functions $U_{pq}$, such that $U_{pq}(z,p,(s_{p},s_{q}))$ denotes player $p$’s utility in a two-player game against player $q$. In standard polymatrix games, a player $p$’s utility is the sum over all of their individual two-player games:
@@ 2 Polynomial Expectation Property for Max-Variant Polymatrix Games
By construction, $M$ is nonincreasing during the loop, so it suffices to subtract $x_{q}(k)$ from $c_{q}$ once $u_{q}(k)=M$ (i.e. $u_{q}(k)$ becomes the “current" utility). Note that, in the case of ties, we can assign the same (arbitrary) tiebreaking rule to the algorithm’s “argmax" and to the “>>" operator, and this invariant will remain true.
@@
  $E(i)\leftarrow 0$ $\triangleright$ exp. utility given fixed action $i$
  $c_{q}\leftarrow 1,\,\,\forall q\neq p$ $\triangleright$ “CDF" variable for each player
@@
   $E(i)\leftarrow E(i)+u_{q^{*}}(0)\cdot(x_{q^{*}}(0)\cdot\prod_{q\neq q^{*}}c_{q})$
@@
   Remove action 0 from $u_{q^{*}}$, shift other actions up
@@ 3.2 Necessity of Polynomial Expectation
Part of the spirit of the original question proposed by Papadimitriou and Roughgarden was to address cases of games without polynomial expectation property, and argue whether they still have polynomial correlated equilibrium schemes. As it turns out, the example they provided, Max-Polymatrix, does indeed have the polynomial expectation property. This inspires the following general question: Papadimitriou and Roughgarden demonstrate that the polynomial expectation property is a sufficient condition for a polynomial correlated equilibrium scheme. To what extent is it necessary?
A small idea in this direction relates to the necessity of the computation of $UX^{\top}$ in the original algorithm in [1]. In particular, to solve the dual program $UX^{\top}]\alpha\geq 0$, it is not entirely necessary to compute $UX^{\top}$ itself; it suffices to have a separation oracle which, given some input $\alpha$, returns either that it is feasible or a halfspace it violates (via the ellipsoid algorithm). This is equivalent to having an oracle which, given a distribution $X^{\top}\alpha$ [^1] Note that the entire vector $X^{\top}\alpha$ might be hard to compute, as we are linearly combining very large vectors. However, we can keep the representation as a linear combination of product distributions and treat “accessing” any element in this distribution as taking polynomial-time instead of constant., returns either that it is a correlated equilibrium or demonstrates a failed constraint. So, we have:
$\text{Verifying CE in poly-time}\implies\text{Finding CE in poly-time}$